# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v6

@@ 8 Container Density
Now that we have established interior regularity and interior degree doubling, we can start to think about what those things mean. The Weak Decreasing Exteriors Lemma implies that containers will decrease in size. We need to see how that decrease in size will impact the degrees of the nodes.
@@
Regular Exterior Degrees
Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree-doubling node. Let a minimum degree node of $G$ be $v_{0}$, and suppose every node in $G$ has the Decreasing Sequence Property (DSP). Then the nodes in a container in $G$ will result in regular exterior degrees.
@@
Assume, for contradiction, that the sizes of their exterior sets differ by at least one. Without loss of generality, let there exist a node $y$ such that $|ext(x_{i},y)|=|ext(x_{j},y)|+1$. This implies that the exterior degree (the number of nodes at distance $i+1$ from $v_{0}$) is irregular between $x_{i}$ and $x_{j}$.
By Lemmas 7.2 (Interior Degree Doubling) and 6.5 (No Back Arcs), we can imply Lemma 5.6 (Weak Decreasing Exteriors Lemma) here. This says that with each distance from the minimum degree node, our exterior size is decreased by 1.
This stopped at the container $C_{i}$. That means there will be two neighboring containers of equal size, $C_{i-1}$ and $C_{i}$. So, a node in $C-{i-2}$ can have its exterior degree double. By 7.2 and 7.1 (Indistinguishably of Regular Nodes), we know that this node has its interior degree double as well.
So, this assumption led to a contradiction.
This is an important property because it ties into the Decreasing Sequence Property (DSP). We were able to use the Weak Decreasing Exteriors Lemma in this proof because the two go hand in hand.
Anything that leads to a hiccup in the decrease in exteriors would cause a node’s degree to double. The DSP is a bottleneck in this way because it is really limiting what can take place in this graph.
@@
Strong Decreasing Exteriors Lemma
Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node in $G$. We are attempting to prevent a degree-doubling node, so for any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of the exterior neighbors must be bounded by
@@
We will prove the statement by induction on the containers $C_{i}$ relative to $v_{0}$.
Base Case ($i=1$):
Consider a first neighbor $x_{1}\in C_{1}$ (the nodes at distance 1 from $v_{0}$). The size of the exterior set
@@
The immediate takeaway from this lemma is the decreasing sequence of exteriors that we were attempting to prove. It is for this sequence that the term ”decreasing sequence property” was originally named.
@@
Interior Degrees Lower Bound
Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node in $G$. Suppose every node in $G$ has the Decreasing Sequence Property (DSP). Let $x$ be a node at distance $i$ from $v_{0}$, and $v_{i-1}$ be a node at distance $i-1$ from $v_{0}$. Then the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$, provided $(v_{0},x)\in G$.
@@
Let $x$ be a node at distance $i$ from $v_{0}$. By Lemma 8.2 (Strong Decreasing Exteriors Lemma), we know that the exterior degree of $v_{i-1}$, $|ext(v_{i-1},x)|$, is bounded by $d^{+}(v_{0})-i$. This places a limit on the number of nodes $x$ can map to at distance $i+1$ from $v_{0}$.
Now, since $x$ has at least the same out-degree as $v_{0}$, $x$ must map to at least $i$ other nodes. Given that back arcs are not asked, $x$ cannot map to nodes at earlier distances.
Therefore, the only remaining possibility is that $x$ maps to interior nodes. Since $x$’s degree must account for at least $i$ mappings. These must be to interior nodes. So, we conclude that $x$ must map to at least $i$ interior nodes.
Thus, $|int(v_{i-1},x)|\geq i$, as required.
Having already placed an upper bound on the exterior sizes and containers, it is only natural that there is an opposite bound for the interiors. Once again, the minimum degree node is the driving force behind this. The degree of these interior nodes will also be impacted by their distance to the minimum degree node.
@@
Regular Graph
All Nodes Have Equal Out-Degree
Let $G$ be an oriented graph represented by a GLOVER. Suppose $v_{0}$ is a minimum degree node of $G$. Further, let all nodes satisfy the Decreasing Sequence Property (DSP). We are attempting to prevent a degree-doubling node. Then for any node $v$ in $G$, the degree of $v$ is equal to $d$, where $\delta=d^{+}(v_{0})$.
@@
With the proofs on the lack of back arcs and the regularity of interior arcs and the regularity of exterior arcs, we arrive at the following equation for each container:
$\delta=|int(v_{p},u_{i})|+|ext(v_{p},u_{i})|$
where $u_{i}$ is in container $C_{i}$, and $v_{p}$ is a parent of $u_{i}$ in container $C_{i-1}$.
@@
This more rapid reduction in the interior degree leads to container density faster. To avoid this, regularity is key.
Thus, every node in every container’s degree is set by the minimum degree node. This leads to the regularity of the graph.
@@ 9 Transitive Triangles and Seymour Diamonds
Before we continue with the development of our DSP algorithm and show how it solves the Seymour Second Neighborhood Conjecture, we need to address some potential problems and show that they will not hinder our algorithm.
@@
Here, we have $x$ pointing to $y$ and $u$ and $y$ pointing to $u$. The concept of a transitive triangle is an important one because it involves the risk of double counting. The node $u$ is both a first neighbor of $x$ and a second neighbor of $x$, so if we do not properly deal with transitive triangles, we might improperly count out second neighbors and thus not have a correct algorithm.
@@
Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, and $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
The Seymour Diamond is a graph theory concept in graph theory where a node $x$ points to two out-neighbors $y$ and $w$, and these two out-neighbors both share the same out-neighbor, $u$. This is a four-vertex-oriented graph. What this situation presents is redundant second neighbors. There is a path $x\rightarrow u_{1}\rightarrow u_{2}$ in this diamond, and there is a path $x\rightarrow u_{2}\rightarrow u_{1}$ in this diamond. As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and, more importantly, the Seymour’s Second Neighborhood Conjecture.
@@
Transitive Triangles from Containers
Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP). For any two consecutive containers $C_{i}$ and $C_{i+1}$, there exist two parent nodes $u\in C_{i}$ and $v\in C_{i}$, and a shared child node $w\in C_{i+1}$, such that the nodes $u$, $v$, and $w$ form a transitive triangle.
Lemma 9.1 (Transitive Triangles from Containers) informs us that transitive triangles are intricate parts of neighboring containers. As such, they do not hinder the algorithm.
@@
The following arcs exist:
1. There are forward arcs $x\to y$ and $y\to u$,
2. There are forward arcs $x\to w$ and $w\to u$,
3. No back arcs exist between containers.
@@ 2. There are forward arcs $x\to w$ and $w\to u$,
Seymour Diamonds and Decreasing Path Algorithm
Suppose $G$ is an oriented graph represented by the GLOVER, where all nodes have the Decreasing Sequence Property (DSP). We are attempting to prevent a degree-doubling node, where we are applying the DSP algorithm. Suppose $G$ contains a subgraph $D$ that forms a Seymour diamond. Then, the presence of the Seymour diamond $D$ does not disrupt the DSP algorithm.
@@
Suppose $G$ is an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node with degree $d$. Consider a transitive triangle formed by nodes $z$, $y$, and $x$, where $d^{+}(y)=1$ and $d^{+}(z)=0$.
@@
  • $d^{+}(y)=d$ and $d^{+}g(z)=d$
  • $d^{+}(y)=d$ and $d^{+}(z)>d$
  • $d^{+}(y)>d$
  • $d^{+}(y)<d$
@@
  • Case 1: $d^{+}(y)=d$ and $d^{+}(z)=d$:
    Both $y$ and $z$ have exactly $d$ outgoing connections. For the triangle to be transitive, they must not share all neighbors. This means that there exists at least one neighbor of $y$ that is not a neighbor of $z$. Thus, $N^{+}(y)\nsubseteq N^{++}(z)$.
  • Case 2: $d^{+}(y)=d$ and $d^{+}(z)>d$:
    Then $z$ may cover some neighbors of $N^{+}(y)$. Since $y$ must map to at least one neighbor unique from those of $z$, it follows that $y$ has a neighbor that $z$ does not connect to.
  • Case 3: If $d^{+}(y)>d$:
    $y$ has more outgoing connections than required. It can utilize these additional connections to map to unique neighbors without impacting the minimum degree condition, ensuring $y$ retains neighbors not covered by $z$.
  • Case 4: If $d^{+}(y)<d$:
    This scenario contradicts the minimum degree requirement since each node must have at least $d$ outgoing connections. Therefore, $y$ cannot be part of the triangle involving $v_{0}$.
Now that we see that transitive triangles and Seymour diamonds are not a problem for our algorithm, we can continue our investigation of oriented graphs. Next, we will blend everything together into our main results.
@@ 10.1 Introduction
Here, we will sum up all the lemmas we have derived from our analysis of oriented graphs. We will state the final algorithm and some final results.
So far, this battle has been a battle between a catalyst, the chosen minimum degree node and the bottleneck, the Decreasing Sequence Property (DSP). We have proved some important things: Back arcs cannot exist; interior degrees must double inside each container, and the exterior degrees decrease in size; interior degrees proportionally increase in size.

### v7

@@ 8 Container Density
Now that we have established interior regularity and interior degree doubling, we can start to think about what those things mean. The Weak Decreasing Exteriors Lemma implies that containers will decrease in size. We need to see how that decrease will impact the degrees of the nodes, the other properties of containers, and the development of our algorithm.
@@
Regular Exterior Degrees Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree-doubling node. Let a minimum degree node of $G$ be $v_{0}$, and suppose every node in a container $C_{i}$ has the Decreasing Sequence Property (DSP). Then the nodes in $C_{i}$ will result in regular exterior degrees.
@@
Assume, for contradiction, that the sizes of their exterior sets differ by at least one. Without the loss of generality, let there exist a node $y$ such that $|ext(x_{i},y)|\geq|ext(x_{j},y)|+1$. This implies that the exterior degree (the number of nodes at distance $i+1$ from $v_{0}$) is irregular between $x_{i}$ and $x_{j}$.
Because Lemmas 7.2 (Interior Degree Doubling) and 6.5 (No Back Arcs) both hold, we can apply Lemma 5.6 (Weak Decreasing Exteriors Lemma) here. This says that with each increase in distance from the minimum degree node, our exterior size is decreased by 1.
This decreasing of the exteriors (and thus the containers) is stopped at the container $C_{i}$. This is where our assumption of the different exterior sizes comes into play. That means there will be two neighboring containers of equal size, $C_{i-1}$ and $C_{i}$. So, a node in $C-{i-2}$ can have its exterior degree double. By 7.2 and 7.1 (Indistinguishably of Regular Nodes), we know that this node has its interior degree double as well.
So, this assumption led to a contradiction. ∎
This is an important property because it ties into the Decreasing Sequence Property (DSP). We were able to use the 5.6 (Weak Decreasing Exteriors Lemma) in this proof because the two go hand in hand.
Anything that leads to a hiccup in the decrease in exteriors would cause a node’s degree to double. The DSP is a bottleneck in this way because it is really limiting what can happen in this graph in the instance of not having a node’s degree double.
@@
Strong Decreasing Exteriors Lemma Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node in $G$. We are attempting to prevent a degree-doubling node. For any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of the exterior neighbors, and thus $C_{i}$ must be bounded by
@@
We will prove the statement by induction on the containers $C_{i}$ relative to $v_{0}$. Base Case ($i=1$): Consider a first neighbor $x_{1}\in C_{1}$ (the nodes at distance 1 from $v_{0}$). The size of the exterior set
@@
In Section 5 (Theoretical Foundation), we demonstrated that mathematical induction could be applied to our framework. This proof highlights the utility of the lemma. It allowed us to construct a proof based on the container distance from the chosen minimum-degree node. It is possible that this result could have been proven without relying on induction. We could have used a different approach or simply avoided induction based on container distance from the minimum degree node entirely. Still, the use of induction provides an important method, particularly for graphs where the usage of induction is more limited. Induction may not always be necessary, but having a variety of alternative approaches to proofs strengthens our theoretical foundation and enhances our flexibility in tackling these problems.
@@
Interior Degrees Lower Bound Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node in $G$. Suppose every node in $G$ has the Decreasing Sequence Property (DSP). Let $x$ be a node at distance $i$ from $v_{0}$, and $v_{i-1}$ be a parent of $x$. Then the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$.
@@
Let $x$ be a node at distance $i$ from $v_{0}$.
By Lemma 8.2 (Strong Decreasing Exteriors Lemma), we know that the exterior degree of $v_{i-1}$, $|ext(v_{i-1},x)|$, is bounded by $d^{+}(v_{0})-i$. This places a limit on the number of nodes $x$ can map to at distance $i+1$ from $v_{0}$.
Now, since $x$ has at least the same out-degree as $v_{0}$, $x$ must map to at least $i$ other nodes. Given that back arcs are not allowed, $x$ cannot map to nodes at earlier distances.
Therefore, the only remaining possibility is that $x$ maps to interior nodes. These must be to interior nodes. So, we conclude that $x$ must map to at least $i$ interior nodes.
Thus, $|int(v_{i-1},x)|\geq i$, as required. ∎
Having already placed an upper bound on the exterior sizes and containers, it is only natural that there is a contrasting bound for the interiors. Once again, the minimum degree node is the driving force, the catalyst, behind this. The degree of these interior nodes will also be impacted by their distance to the minimum degree node.
@@
Regular Graph All Nodes Have Equal Out-Degree Let $G$ be an oriented graph represented by a GLOVER. Suppose $v_{0}$ is a minimum degree node of $G$. Further, let all nodes satisfy the Decreasing Sequence Property (DSP). We are attempting to prevent a degree-doubling node. Then for any node $v$ in $G$, the degree of $v$ is equal to $d$, where $\delta=d^{+}(v_{0})$.
@@
This more rapid reduction in the interior degree leads to container density faster, which would simply lead to a degree doubling node faster. To avoid this, overall regularity is key.
Thus, every node in every container’s degree is set by the minimum degree node. This leads to the regularity of the graph. ∎
@@
Example 8.1.
0
$\Rightarrow$ 111122223333444455556666777788889999101010101111111112121212131313131414141415151515161616161717171718181818191919192020202021212121222222222323232324242424H
[…]
Example 8.1 shows some of the patterns that emerge with this regularity in the SSNC. Starting from the outside in, going inward represents the order of containers as they increase in distance from $v_{0}$. This example has visualized only the interior arcs, as the exterior arcs would complicate this visualization. However, just as in previous examples, all nodes in a previous container are mapping to all nodes in the next container.
Some of those patterns that emerge are first the equal out-degree of nodes. For even-sized containers, we are able to see symmetry about balance, uniformity, and a sense of order within each container. Also, while it is not a point of the conjecture, we do see the presence of cycles in each container.
@@ 9 Transitive Triangles and Seymour Diamonds
Before we continue with the development of our algorithm and show how it solves the SSNC, we need to address some potential problems and show that they will not hinder our algorithm.
@@
Here, we have $x$ pointing to $y$ and $u$ and $y$ pointing to $u$. The concept of a transitive triangle is an important one because it involves the risk of double counting. The node $u$ is both a first neighbor of $x$ and a second neighbor of $x$. If we do not properly deal with transitive triangles, we might improperly count out second neighbors and thus not have a correct algorithm.
@@
Let $x\in G$ and $y,w\in N^{+}(x)$. Then $x$, $y$, $w$, and $u$ form a Seymour diamond if $(y,u)$ and $(w,u)\in G$.
The Seymour Diamond is a graph theory concept in graph theory where a node $x$ points to two out-neighbors $y$ and $w$, and these two out-neighbors both share the same out-neighbor, $u$. This is a four-vertex-oriented graph. What this situation presents is redundant second neighbors. There is a path $x\rightarrow y\rightarrow u$ in this diamond, and there is a path $x\rightarrow w\rightarrow u$ in this diamond. As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and, more importantly, the Seymour’s Second Neighborhood Conjecture.
@@
Figure 12: Illustration of a transitive triangle with parent node $x$, child nodes $y$ and $u$, where $u$ is also a child of $y$
@@
Transitive Triangles from Containers Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP). For any two consecutive containers $C_{i}$ and $C_{i+1}$, there exist two parent nodes $u\in C_{i}$ and $v\in C_{i}$, and a shared child node $w\in C_{i+1}$, such that the nodes $u$, $v$, and $w$ form a transitive triangle.
Lemma 9.1 (Transitive Triangles from Containers) informs us that transitive triangles are intricate parts of neighboring containers. As such, they do not hinder the algorithm. Transitive triangles represent a common situation with our containers. A parent node $x$ has two children $y$ and $u$ that are also connected by an arc. The purpose of the interior arcs is to ensure that the parent node $x$’s degree does not double. We have already shown through 8.3 (Interior Degrees Lower Bound) that every node in Container $i$ has an interior degree at least $i$.
The two arcs from a parent node and the interior nodes are what create these transitive triangles. What is not acknowledged in lemma 9.1 (Transitive Triangles from Containers) is that every node in the container takes both non-parent roles in different transitive triangles. Every node is both a child of a parent node in the previous container and has a degree of at least one. This is what creates the transitive triangles in every container.
@@
The following arcs exist: 1. There are arcs $x\to y$ and $y\to u$, 2. There are arcs $x\to w$ and $w\to u$, 3. No back arcs exist between containers.
@@
This is exactly what we are seeing happen every time there is a Seymour diamond present between three containers. There is a parent node, a choice of at least 2 interior nodes, and an exterior node. Not only is this not a problem; it’s one of the benefits of the GLOVER data structure. Those nodes $y$ and $w$ represent equidistant nodes from $v_{0}$ whose job is to ensure that $x$’s degree does not double. The arcs from $y$ to $u$ and $w$ to $u$ are exterior arcs, connecting these nodes to the next container. The purpose of these arcs is first to meet the degree requirements of each node and second to ensure connectivity. Once each node $y$ and $w$ has connected to the third container, it can take a role as a parent and form its own Seymour diamond.
@@
Figure 13: Illustration of a Seymour diamond with parent node $x$, child nodes $y$ and $w$, and grandchild node $u$
@@
Seymour Diamonds and Decreasing Path Algorithm Suppose $G$ is an oriented graph represented by the GLOVER, where all nodes have the Decreasing Sequence Property (DSP). We are attempting to prevent a degree-doubling node, where we are applying the DSP algorithm. Suppose $G$ contains a subgraph $D$ that forms a Seymour diamond. Then, the presence of the Seymour diamond $D$ does not disrupt the DSP algorithm.
@@
Lemma 9.3 (Seymour Diamonds and Decreasing Path Algorithm) shows how we have been proceeding through the GLOVER data structure through our Seymour diamonds. These diamonds are constructed to be presented in our containers. As a result, we simply select an interior node within the container and continue. This is what we saw in our examples in section 4, where Seymour diamonds were also present, for example 4.3. This is also what we saw in our social media example 5.5 in section 5. This is also what we saw in our regular graph example in 8.1 in section 8.
@@
Suppose $G$ is an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node with degree $\delta$. Consider a transitive triangle formed by nodes $u$, $y$, and $x$, where $d^{+}(y)=1$ and $d^{+}(u)=0$.
@@
  • $d^{+}(y)=\delta$ and $d^{+}(u)=\delta$
  • $d^{+}(y)=\delta$ and $d^{+}(u)>\delta$
  • $d^{+}(y)>\delta$ and $d^{+}(u)=\delta$
  • $d^{+}(y)>\delta$ and $d^{+}(u)>\delta$:
@@
  • Case 1: $d^{+}(y)=\delta$ and $d^{+}(u)=\delta$: Both nodes $y$ and $u$ out-degree $\delta$. The triangle has the arc $(y,u)$. Since the graph is oriented, the arc $(u,y)$ cannot belong to $G$. This means that $u$ has $\delta$ out-neighbors that are not in the triangle. Since we assumed that $d^{+}(y)=\delta$ and $d^{+}(u)=\delta$, we know that $y$ cannot connect to all not $u$’s neighbors since $y$ has a neighbor in the triangle ($u$). This node that is $u$’s neighbor, but not $y$’s neighbor, is a candidate to extend the path.
  • Case 2: $d^{+}(y)=\delta$ and $d^{+}(u)>\delta$: The node $y$ has out-degree $\delta$. The node $u$ has out-degree greater than $\delta$. Again, the triangle has the arc $(y,u)$. Since the graph is oriented, the arc $(u,y)$ cannot belong to $G$. This means that $u$ has more than $\delta$ out-neighbors that are not in the triangle. We assumed that $d^{+}(y)=\delta$ and $d^{+}(u)>\delta$. We know that $y$ cannot connect to all not $u$’s neighbors since $d^{+}(u)>d^{+}(y)$. We can then select one of these nodes that is a neighbor of $u$, but not a neighbor of $y$ to be a candidate to extend the path.
  • Case 3: If $d^{+}(y)>\delta$ and $d^{+}(u)=\delta$: The node $y$ has out-degree greater than $\delta$. The node $u$ has out-degree $\delta$. The triangle has the arc $(y,u)$. Since the graph is oriented, the arc $(u,y)$ cannot belong to $G$. The node $y$ has more than $\delta$ out-neighbors that are not in the triangle. The node $u$ has exactly $\delta$ out-neighbors that are not in the triangle. The maximum number of neighbors that $y$ and $u$ can share is $\delta$, that being $u$’s out-degree. Then we can select one of the nodes that is a neighbor of $y$, but not a neighbor of $u$ to be a candidate to extend the path.
  • Case 4: If $d^{+}(y)>\delta$ and $d^{+}(u)>\delta$: The node $y$ has out-degree greater than $\delta$. The node $u$ has out-degree greater than $\delta$. We have a few cases because we have to remember that in this data structure the purpose of our containers and the interior nodes was to prevent parent nodes from having their degrees double. The case where both nodes in a transitive triangle have more than $\delta$ out-neighbors means that $x$ has up to $2*\delta$ out-neighbors, which would cause its degree to double. Thus, to prevent $x$’s degree from doubling, there must be some overlap between the out-neighbors of $y$ and the out-neighbors of $u$. These two sets cannot be equal to one another, though. The most overlap we can have is $N^{+}(u)\subsetneq N^{+}(y)$ or $N^{+}(y)\subsetneq N^{+}(u)$. Both of these situations leave room for a node in the larger set that is not in the smaller set as a candidate to extend the path. As we allow for less overlap, the number of candidate nodes to extend the path grows. We can see that there will always be a node in the larger set to extend the path.
Now that we have a better understanding of transitive triangles and Seymour diamonds and how they interact with our data structure. We can see that they are not a problem for the development of our algorithm. We will continue our investigation of oriented graphs. Next, we will blend everything together into our main results.
@@ 10.1 Introduction
Here, we will sum up all the lemmas we have derived from our analysis of oriented graphs. We will state the final algorithm and some results.
So far, this battle has been a battle between a catalyst, the chosen minimum degree node, and the bottleneck, the Decreasing Sequence Property (DSP). We have proved some important things: Back arcs cannot exist; interior degrees must double inside each container, and the exterior degrees decrease in size; interior degrees proportionally increase in size.